OpenStudy (gabylovesyou):

Which of the following is not equivalent to the formula for the circumference of a circle

OpenStudy (gabylovesyou):

C = 2 Pi r

OpenStudy (gabylovesyou):

C/r pi = 2 C/2 pi = r C/2r = pi 2 pi/c = r